# Title of Your Paper

### v1

@@ 2.2 Quantifying Fairness: A Review of Mathematical Fairness Metrics
Despite these limitations, the core methodology outlined in Section 5 is adopted for the analysis within this paper. It introduces the concept of relational evaluation of fairness, acknowledging and operating under the inherent complexities, practical constraints, and ethical dilemmas prevalent in this domain. This approach allows for a more nuanced evaluation that considers crucial factors like financial inclusion alongside the potential negative impacts on underserved groups, moving beyond simplistic metrics. Despite stated limitations, the methodology’s strength lies in its more holistic view of fairness, ability to capture the ethical trade-offs that decision-makers are regularly confronted with, and its encouragement of explicit prioritization of identified values and objectives.
2.4 A Review of ML Algorithms
2.4.1 Logistic regression (LR)
2.4.2 K-nearest neighbors classification (KNN)
The K-nearest neighbors (KNN) algorithm is a non-parametric, supervised machine learning approach utilized for classification and regression tasks. In the KNN algorithm, data points are assigned to a class based on the majority class of their ’K’ nearest neighbors in the feature space. The ’K’ in KNN signifies the number of neighbors considered for the classification or regression decision. In the context of our focus on classification tasks, the algorithm assigns a class label to a data point through a majority vote from its nearest neighbors. For regression tasks, KNN computes the average value of the K nearest neighbors to predict a continuous outcome. KNN is versatile and does not assume any underlying data distribution, making it suitable for various types of datasets. However, its performance can be sensitive to the choice of the distance metric and the value of K, which may become computationally expensive for large datasets. Additionally, it suffers from the curse of dimensionality if the dimensions of the feature space are large, as there will generally be very few training observations close to the point of interest. For a recent review and in-depth analysis of KNN, we refer interested readers to Syriopoulos et al. (2023). For this study, we use K=5 as specified in Lee and Floridi (2021).
2.4.3 Classification and regression tree (CART)
Classification and regression trees (CART) introduced by Breiman et al. (1984) constitutes a nonparametric supervised learning approach that builds tree-like models for both categorical (classification) and continuous (regression) prediction tasks. The algorithm recursively partitions the feature space into distinct regions, assigning a predictive model (either class labels for classification or continuous values for regression) to each resulting segment. CART employs binary splitting at each node, selecting the optimal feature and threshold based on criteria such as Gini impurity for classification or mean squared error for regression. The resulting tree structure provides interpretable decision paths, helping to understand complex relationships within the data. While CART inherently handles non-linear relationships and interactions, it may be prone to overfitting, prompting the adoption of techniques like pruning and ensemble methods for regularization. However, it is an unstable algorithm, particularly when dealing with noisy data or in the presence of outliers.
2.4.4 Gaussian Naïve Bayes (NB)
Gaussian Naïve Bayes (NB) is a probabilistic generative model designed for classification tasks, known for its simplicity and efficiency. It is derived from Bayes’ theorem and assumes that features are conditionally independent given the class label, thereby simplifying the computation of probabilities. In the context of continuous features, Gaussian NB models the likelihood of each feature’s distribution as Gaussian, facilitating the estimation of class probabilities. This algorithm is particularly well-suited for datasets where feature-independence assumptions hold to a reasonable extent. Despite its inherent simplicity and the naive assumption of feature independence, Gaussian NB often performs remarkably well, especially in scenarios with limited training data. A recent review including software implementation and extensions of the method can be found in Wickramasinghe and Kalutarage (2021).
2.4.5 Random Forest (RF)
Random forest (RF) is an ensemble learning approach for both classification and regression tasks, known for harnessing the collective strength of multiple decision trees to deliver robust and well-generalized predictions (Breiman, 2001). It operates by creating bootstrap samples from the training data and randomly selecting features during tree induction. The ensemble’s prediction is aggregated through majority voting (classification) or averaging (regression) to reduce variance and improve model accuracy. This ensemble approach effectively addresses overfitting issues common in single decision trees, while maintaining feature importance measures for understanding the impact of individual features on the prediction. However, choosing the optimal number of trees and tuning hyperparameters remain crucial steps in optimizing RF performance in various applications. Although there are newer techniques for pruning RF (Manzali and Elfar (2023)), we use the hyperparameter specifications from Lee and Floridi (2021).
2.5 Model Multiplicity and Less Discriminatory Algorithms (LDAs)
@@
Given the increasing reliance on AI systems in decision-making processes, it is imperative to carefully select models that do not disproportionately impact marginalized segments of society. The principle of model multiplicity suggests that prediction models may share a similar predictive performance (accuracy), but may differ in other aspects such as fairness, explainability, or robustness. Model multiplicity can be a result of either procedural multiplicity or predictive multiplicity (Dai et al., 2025).
@@
Procedural multiplicity describes how models with different internal processes may come to equally accurate predictions (Dai et al., 2025). Their internal differences may lie, in part, in the weight they give to different inputs or the variables they use altogether. Interestingly, they can reach the same outcome. For instance, a random forest and a linear model might achieve comparable levels of accuracy on a given task yet arrive at their predictions through fundamentally different reasoning processes (Black et al., 2022). Anders et al. (2020) illustrate this point through a study of credit-scoring models: although two models produced identical predictions across all data points, their underlying justifications diverged significantly, one model’s decisions were influenced by gender, while the other relied on variables such as income and tax history.
@@
Predictive multiplicity describes the phenomena in which different models, despite achieving the same overall accuracy on a given task, produce conflicting predictions for specific inputs (Black et al., 2022). Marx et al. (2020) define predictive multiplicity as “the ability of a prediction problem to admit competing models with conflicting predictions”. In other words, two models might perform equally well in aggregate but still assign different (or even contradicting) labels to the same data point. This phenomenon is assessed using the labeled datasets accessible to model developers, typically the training or test sets. As developers only have access to a finite amount of labeled data, they cannot fully observe all potential points of disagreement between models. Instead, they must rely on estimates drawn from the data at hand.
2.5.1 Model Multiplicity and Fairness
@@
Two prominent themes have emerged in the literature on model multiplicity, one focusing on the opportunities for selecting algorithms that confer a greater fairness with little or no cost to accuracy, the other raises concerns about the apparent arbitrariness of prediction (Dai et al., 2025).
@@
Contrary to the fixed model assumption, the phenomenon of multiplicity implies that users can select models which are best aligned with their values such as fairness or interpretability without giving up accurate predictions (Black et al., 2022).
@@
Sokol et al. (2024), while applying large real-life datasets, demonstrates the prevalence of unfairness across models with equal accuracy, and the difficulty to mitigate the effect to grant each individual the most favorable outcome. Black argues that fairness (individual and group) in model predictions is not a necessary trade-off for accuracy in model selection. Building on this idea, Black et al. (2022) proposes that model multiplicity can be interpreted in a way that aligns with the disparate impact doctrine specifically, by encouraging organizations to actively explore a range of equally accurate models in search of those that reduce discriminatory outcomes. These alternative models would need to match the performance of a baseline model already considered suitable for deployment.
@@
Creel and Hellman (2021) touch on the troubling consequences of predictive multiplicity, particularly its potential to introduce arbitrariness into algorithmic decision-making. However, they ultimately contend that this arbitrariness becomes a serious concern only in scenarios where an algorithmic monoculture exists—where the same type of model or system is used so pervasively that individuals are systematically denied opportunities across multiple domains.
@@
This suggests that unless model developers clearly define and optimize specific behaviors like fairness, robustness, or interpretability, there is little reason to expect a model will naturally demonstrate those qualities. This challenge, referred to as the problem of under specification by D’Amour et al. (2020), highlights the importance of explicitly articulating the properties we want in our models. Without doing so, those desired characteristics are unlikely to emerge on their own. Dai et al. (2025) propose that selecting for accuracy alone will lead to a much less fair outcome, and recommend explicitly optimizing for fairness when selecting for an algorithm.
@@
Practically, the theory of model multiplicity raises several questions, particularly regarding fairness and justifiability. While model multiplicity expands the range of options available during model selection, it may simultaneously place responsibility on developers to use that flexibility thoughtfully and to provide clear justifications for their chosen approaches (Black et al., 2022). This responsibility has been referred to as the obligation to search for less discriminatory algorithms (LDAs), outlined in the following section.
2.5.2 Less Discriminatory Algorithms
@@
In recent years, the concept of LDAs has gained traction at the intersection of law, ML fairness, and compliance. LDAs are, as the name suggests, algorithms that can achieve a specified task with notably less disproportionate harm or negative consequences to individuals in marginalized groups. A growing body of literature has explored the LDA paradigm which Black et al. (2024a) proposed, particularly within high-impact algorithmic contexts, as well as the technical and practical constraints of defining and operationalizing such searches.
@@
Black et al. (2024b) and Gillis et al. (2024) lay the legal and normative foundations for LDA obligations, with Black et al. (2024b) cleanly articulating a case for a duty to search for LDAs as a necessary component of algorithmic accountability to avoid liability under the existing anti-discrimination legal regime.Gillis et al. (2024) extends this concept through addressing the challenge of operationalizing such a search, arguing for practical tools and guidelines that help organizations move from a legal requirement to technical implementation. Their approach focuses on treating the search as an optimization problem, both identifying linear classification models which minimize discrimination given a set of business constraints, and return cases where this was not possible within the originally stated criteria, offering conclusive results either way.
@@
Black et al. (2023) previously called specifically for a pipeline aware research agenda, urging fairness scholars to develop concrete tools for evaluating discrimination across end-to-end model pipelines, rather focusing solely on a model’s outputs. Building on this directive, Laufer et al. (2025) continues to raise critical questions around what qualifies as a “less discriminatory algorithm”, arguing that this concept must be conceptualized in terms of competing objectives such as utility, risk tolerance, and business constraints. Laufer et al. (2025) also directly addresses the inherent limits of LDA search, noting that the space of less discriminatory models is vast and often underspecified, and proposes ways to bound this space through reasonable temporal and resource constraints, optimization principles, and fairness constraints. Helpfully, they also demonstrate that while businesses and consumers may have different utility priorities, there is room for consideration of both if included as relevant variables from the outset of model development.
@@
Across these diverse literatures, fairness work has greatly advanced, seen through a recent spur of measurement and evaluation methods. While the legal field has articulated the normative obligation to identify LDAs, there is still significant work to be done to convert this into practice in ways that are transparent, replicable, and lightweight. Existing LDA search evaluation methods, by and large, test for fairness extensively within a single model family, relying on technically demanding processes including complex hyperparameter optimization routines, which may require non-insignificant computational resources, or technical expertise. We refer to this approach in this paper as a ’vertical LDA search’. Although rigorous and useful, relying exclusively on this approach poses practical challenges for real-world implementation, particularly in resource-constrained environments and reporting contexts, which may pose barriers to industry-scale adoption of LDAs. Our contribution seeks to bridge these gaps to encourage LDA usage as widely as possible, building on the research agenda outlined by Black et al. (2023), particularly the urgent need for operational tools and methods in high-risk industries.
@@
3 Methods and Methodology
This section outlines the empirical design and methodological approach used to explore the fairness-performance trade-offs we chose across ML model families in mortgage lending. We begin by detailing the primary dataset used within this study, the 2021 Home Mortgage Disclosure Act (HMDA) data, published by the Consumer Financial Protection Bureau (2021), which extensively documents records on mortgage lending practices in the United States, providing the foundation for analysis. The approach to model development for our five algorithms, including our target variable and relevant feature encoding, is disclosed. Finally, we outline the methodological assumptions underpinning our analysis, including the choices made to simulate real-world lending conditions, which attempt to balance clarity and reproducibility.
3.1 Case Study: HMDA 2021 Dataset
The HMDA dataset categorizes race and ethnicity directly using the applicant_race_1 and applicant_ethnicity_1 fields, as shown in Table 3.1.1. For data exploration purposes, outcomes for Asian, Black, Hispanic, Native American and Pacific Islander, and Non-Hispanic White groups are displayed in Section 3.1, however for graphs included in Section 4.2, only Black and Non-Hispanic White data are used.
Table 3.1.1: Data Fields (Race and Ethnicity) with Values and Definitions
@@ 3.1.1 Data Visualizations
Figure 1: HMDA 2021 Home Ownership Race Distribution
3.2 Model Design and Development
Table 3.2.2 outlines the target variable used in our machine learning model, indicating whether a loan application was approved or denied based on various loan actions. It summarizes the relationship between different loan actions and the resulting approval or denial status, serving as a reference for building predictive models in the loan approval process. In this case, the positive label for prediction is ”Loan Denied,” and the negative label is ”Loan Approved.” For further information on features used to develop the model, reference Table B.1.1. The features provide valuable insight into the profiles of loan applicants and their neighborhoods, which may contribute to the likelihood of loan approval or denial. For data sampling, this paper used the same data sampling technique as stated in Lee and Floridi (2021).
Table 3.2.2: Mapping of Loan Action taken to Approved and Denied
@@ 3.3 Methodological Assumptions
The assumptions used to estimate the impact of a model on financial inclusion and protected groups are as discussed in Lee and Floridi (2021), with the addition of Assumption 7:
@@ 1. Assumption 1
    Data Representativeness: HMDA data are assumed to reflect loan outcomes of a perfectly informed market that relies only on a single loan approval model. Under this assumption, all approved loans would get repaid and all denied loans ultimately defaulted. This aligns with the regulatory framework established by laws like the Fair Credit Reporting Act (FCRA), which requires data used in credit decisions to meet standards of maximum possible accuracy. Thus, the data used is assumed to reflect and inform accurate underwriting practices (CFPB, 2022).
@@ 3. Assumption 3
    1. 1.
      The lender uses one of the models discussed in section 2.4 one step at a time to predict whether the loan will default, i.e., whether or not the mortgage loan application should be approved or denied as discussed in Table 3.2.2.
    2. 2.
@@ 5. Assumption 5
      $\text{EV}=(P\times R)-((1-P)\times L),\text{where}$ (1)
@@ 7. Assumption 7
    Racial Negative Impact: We focused on Black mortgage applicants as a benchmark for potential disparities relative to Non-Hispanic White applicants, reflecting the historical context of housing discrimination. Our metric measured the direct denial rate experienced by Black applicants (denied loans / total Black applicants). While not using Adverse Impact Ratios (AIR) (see Relman Colfax PLLC (2020)), this approach facilitates transparent comparison with Lee and Floridi (2021)’s findings, requiring careful interpretation while ensuring compliance with fair lending principles.
Assumptions 1 to 7 provide the basis for estimating financial inclusion with the expected value of the loans when predicted risks of default are known. The primary objective of financial inclusion is to bridge the gap between those who have access to banking and financial services, including access to mortgage loans, and those who are excluded, often due to poverty or protected characteristics such as race and color, to help reduce inequality and advance economic mobility (Relman Colfax PLLC, 2020). Future work will be needed to investigate the updates required to apply the assumptions to low-income mortgage applicants and high-risk applicants with and without consideration of protected characteristics.
4 Results
In this section, we present the findings of applying Lee and Floridi (2021)’s fairness assessment methodology to contemporary lending data from the 2021 HMDA. Similarly, our analysis evaluates the performance of multiple machine learning models across different fairness metrics, quantifying the trade-offs between accuracy and fairness through variation in algorithmic structure. Overall, our findings demonstrate the robustness of the original framework and its ability to transparently visualize algorithmic trade-offs structured under complex real-world conditions. These conditions are well illustrated through the highly regulated environment that the lending case study presents.
4.1 Fairness Incompatibility and Trade-off Analysis
Racial disparity measures such as those discussed in Lee and Floridi (2021), do not account for historical bias that is prevalent in mortgage data, although they are often used in evaluating lending decisions. For example, existing algorithmic fairness metrics either measure disparity in lending decisions at group or individual levels under absolute mathematical conditions as previously discussed in Section 2.2. Hence, we tested the compatibility of different mathematical notions of fairness (see Section 2 and Figure 2 in Lee and Floridi (2021)) to decide if trade-offs such as those suggested in studies such as the aforementioned, Kleinberg et al. (2016), and Hellman (2020) are necessary for the models in this work.

### v2

@@ 2.2. Quantifying Fairness: A Review of Mathematical Fairness Metrics
Despite these limitations, the core methodology outlined in Section 5 is adopted for the analysis within this paper. It introduces the concept of relational evaluation of fairness, acknowledging and operating under the inherent complexities, practical constraints, and ethical dilemmas prevalent in this domain. This approach allows for a more nuanced evaluation that considers crucial factors like financial inclusion alongside the potential negative impacts on underserved groups, moving beyond simplistic metrics. Despite stated limitations, the methodology’s strength lies in its more holistic view of fairness, ability to capture the ethical trade-offs that decision-makers are regularly confronted with, and its encouragement of explicit prioritization of identified values and objectives. For a closer review of the algorithms employed, see Appendix A.
2.4. Model Multiplicity and Less Discriminatory Algorithms (LDAs)
@@
Given the increasing reliance on AI systems in decision-making processes, it is imperative to carefully select models that do not disproportionately impact marginalized segments of society. The principle of model multiplicity suggests that prediction models may share a similar predictive performance (accuracy), but may differ in other aspects such as fairness, explainability, or robustness. Model multiplicity can be a result of either procedural multiplicity or predictive multiplicity (dai2025).
@@
Procedural multiplicity describes how models with different internal processes may come to equally accurate predictions (dai2025). Their internal differences may lie, in part, in the weight they give to different inputs or the variables they use altogether. Interestingly, they can reach the same outcome. For instance, a random forest and a linear model might achieve comparable levels of accuracy on a given task yet arrive at their predictions through fundamentally different reasoning processes (black2022). anders2020 illustrate this point through a study of credit-scoring models: although two models produced identical predictions across all data points, their underlying justifications diverged significantly, one model’s decisions were influenced by gender, while the other relied on variables such as income and tax history.
@@
Predictive multiplicity describes the phenomena in which different models, despite achieving the same overall accuracy on a given task, produce conflicting predictions for specific inputs (black2022). marx2020 define predictive multiplicity as “the ability of a prediction problem to admit competing models with conflicting predictions”. In other words, two models might perform equally well in aggregate but still assign different (or even contradicting) labels to the same data point. This phenomenon is assessed using the labeled datasets accessible to model developers, typically the training or test sets. As developers only have access to a finite amount of labeled data, they cannot fully observe all potential points of disagreement between models. Instead, they must rely on estimates drawn from the data at hand.
2.4.1. Model Multiplicity and Fairness
@@
Two prominent themes have emerged in the literature on model multiplicity, one focusing on the opportunities for selecting algorithms that confer a greater fairness with little or no cost to accuracy, the other raises concerns about the apparent arbitrariness of prediction (dai2025).
@@
Contrary to the fixed model assumption, the phenomenon of multiplicity implies that users can select models which are best aligned with their values such as fairness or interpretability without giving up accurate predictions (black2022).
@@
sokol2024, while applying large real-life datasets, demonstrates the prevalence of unfairness across models with equal accuracy, and the difficulty to mitigate the effect to grant each individual the most favorable outcome. Black argues that fairness (individual and group) in model predictions is not a necessary trade-off for accuracy in model selection. Building on this idea, black2022 proposes that model multiplicity can be interpreted in a way that aligns with the disparate impact doctrine specifically, by encouraging organizations to actively explore a range of equally accurate models in search of those that reduce discriminatory outcomes. These alternative models would need to match the performance of a baseline model already considered suitable for deployment.
@@
creel2021 touch on the troubling consequences of predictive multiplicity, particularly its potential to introduce arbitrariness into algorithmic decision-making. However, they ultimately contend that this arbitrariness becomes a serious concern only in scenarios where an algorithmic monoculture exists—where the same type of model or system is used so pervasively that individuals are systematically denied opportunities across multiple domains.
@@
This suggests that unless model developers clearly define and optimize specific behaviors like fairness, robustness, or interpretability, there is little reason to expect a model will naturally demonstrate those qualities. This challenge, referred to as the problem of under specification by damour2020, highlights the importance of explicitly articulating the properties we want in our models. Without doing so, those desired characteristics are unlikely to emerge on their own. dai2025 propose that selecting for accuracy alone will lead to a much less fair outcome, and recommend explicitly optimizing for fairness when selecting for an algorithm.
@@
Practically, the theory of model multiplicity raises several questions, particularly regarding fairness and justifiability. While model multiplicity expands the range of options available during model selection, it may simultaneously place responsibility on developers to use that flexibility thoughtfully and to provide clear justifications for their chosen approaches (black2022). This responsibility has been referred to as the obligation to search for less discriminatory algorithms (LDAs), outlined in the following section.
2.4.2. Less Discriminatory Algorithms
@@
In recent years, the concept of LDAs has gained traction at the intersection of law, ML fairness, and compliance. LDAs are, as the name suggests, algorithms that can achieve a specified task with notably less disproportionate harm or negative consequences to individuals in marginalized groups. A growing body of literature has explored the LDA paradigm which Black2023Algorithms proposed, particularly within high-impact algorithmic contexts, as well as the technical and practical constraints of defining and operationalizing such searches.
@@
black2024 and gillis2024 lay the legal and normative foundations for LDA obligations, with black2024 cleanly articulating a case for a duty to search for LDAs as a necessary component of algorithmic accountability to avoid liability under the existing anti-discrimination legal regime. gillis2024 extends this concept through addressing the challenge of operationalizing such a search, arguing for practical tools and guidelines that help organizations move from a legal requirement to technical implementation. Their approach focuses on treating the search as an optimization problem, both identifying linear classification models which minimize discrimination given a set of business constraints, and return cases where this was not possible within the originally stated criteria, offering conclusive results either way.
@@
black2023 previously raised the idea of pipeline-aware fairness research, urging scholars to develop concrete tools for evaluating discrimination across end-to-end model pipelines rather than focusing solely on a model’s outputs. Building on this direction, laufer2025 continues to raise critical questions around what qualifies as a ”less discriminatory algorithm”, arguing that this concept must be conceptualized in terms of competing objectives such as utility, risk tolerance, and business constraints. laufer2025 also directly addresses the inherent limits of LDA search, noting that the space of less discriminatory models is vast and often underspecified, and proposes ways to bound this space through reasonable temporal and resource constraints, optimization principles, and fairness constraints. Helpfully, they demonstrate that while businesses and consumers may have different utility priorities, there is room for consideration of both if included as relevant variables from the outset of model development.
@@
Across these diverse literatures fairness work has greatly advanced, seen through a recent spur of measurement and evaluation methods. The legal field has articulated the normative obligation to identify LDAs, but there is still significant work to be done to convert this into practice in ways that are transparent, replicable, and lightweight. Existing LDA search evaluation methods, by and large, test for fairness extensively within a single model family, relying on technically demanding processes including complex hyperparameter optimization routines, which may require non-insignificant computational resources, or technical expertise. We refer to this approach in this paper as a ’vertical LDA search’. Although rigorous and useful, relying exclusively on this approach poses practical challenges for real-world implementation, particularly in resource-constrained environments and reporting contexts, which may pose barriers to industry-scale adoption of LDAs. Our contribution seeks to bridge these gaps to encourage LDA usage as widely as possible, building on the research agenda outlined by black2023, particularly the urgent need for operational tools and methods in high-impact industries.
@@
3. Methods and Methodology
This section outlines the empirical design and methodological approach used to explore the fairness-performance trade-offs we chose across ML model families in mortgage lending. We begin by detailing the primary dataset used within this study, the 2021 Home Mortgage Disclosure Act (HMDA) data, published by the ffiec_hmda_2021, which extensively documents records on mortgage lending practices in the United States, providing the foundation for analysis. The approach to model development for our five algorithms, including our target variable and relevant feature encoding, is disclosed. Finally, we outline the methodological assumptions underpinning our analysis, including the choices made to simulate real-world lending conditions, which attempt to balance clarity and reproducibility.
3.1. Case Study: HMDA 2021 Dataset
The HMDA dataset categorizes race and ethnicity directly using the applicant_race_1 and applicant_ethnicity_1 fields, as shown in Table 3.1.1. For data exploration purposes, outcomes for Asian, Black, Hispanic, Native American and Pacific Islander, and Non-Hispanic White groups are displayed in 3.1.1 , however for graphs included in4.2, only Black and Non-Hispanic White data are used.
Table 3.1.1. Data Fields (Race and Ethnicity) with Values and Definitions
@@ 3.1.1. Data Visualizations
Figure 1. HMDA 2021 Home Ownership Race Distribution
3.2. Model Design and Development
Table 3.2.2 outlines the target variable used in our machine learning model, indicating whether a loan application was approved or denied based on various loan actions. It summarizes the relationship between different loan actions and the resulting approval or denial status, serving as a reference for building predictive models in the loan approval process. In this case, the positive label for prediction is ”Loan Denied,” and the negative label is ”Loan Approved.” For further information on features used to develop the model, reference Table D.0.1. The features provide valuable insight into the profiles of loan applicants and their neighborhoods, which may contribute to the likelihood of loan approval or denial. For data sampling, this paper used the same data sampling technique as stated in lee21.
Table 3.2.2. Mapping of Loan Action taken to Approved and Denied
@@ 3.3. Methodological Assumptions
The assumptions used to estimate the impact of a model on financial inclusion and protected groups are as discussed in lee21, with the addition of Assumption 7:
@@ 1. Assumption 1
    Data Representativeness: HMDA data are assumed to reflect loan outcomes of a perfectly informed market that relies only on a single loan approval model. Under this assumption, all approved loans would get repaid and all denied loans ultimately defaulted. This aligns with the regulatory framework established by laws like the Fair Credit Reporting Act (FCRA), which requires data used in credit decisions to meet standards of maximum possible accuracy. Thus, the data used is assumed to reflect and inform accurate underwriting practices (cfpb2022accuracy).
@@ 3. Assumption 3
    1. (1)
      The lender uses one of the models discussed in section A one step at a time to predict whether the loan will default, i.e., whether or not the mortgage loan application should be approved or denied as discussed in Table 3.2.2.
    2. (2)
@@ 5. Assumption 5
      (1) $\text{EV}=(P\times R)-((1-P)\times L),\text{where}$
@@ 7. Assumption 7
    Racial Negative Impact: We focused on Black mortgage applicants as a benchmark for potential disparities relative to Non-Hispanic White applicants, reflecting the historical context of housing discrimination. Our metric measured the direct denial rate experienced by Black applicants (denied loans / total Black applicants). While not using Adverse Impact Ratios (AIR) (see relmancolfax_upstart_2020), this approach facilitates transparent comparison with lee21’s findings, requiring careful interpretation while ensuring compliance with fair lending principles.
Assumptions 1 to 7 provide the basis for estimating financial inclusion with the expected value of the loans when predicted risks of default are known. The primary objective of financial inclusion is to bridge the gap between those who have access to banking and financial services, including access to mortgage loans, and those who are excluded, often due to poverty or protected characteristics such as race and color, to help reduce inequality and advance economic mobility (relmancolfax_upstart_2020). Future work will be needed to investigate the updates required to apply the assumptions to low-income mortgage applicants and high-risk applicants with and without consideration of protected characteristics.
4. Results
In this section, we present the findings of applying lee21’s fairness assessment methodology to contemporary lending data from the 2021 HMDA. Similarly, our analysis evaluates the performance of multiple machine learning models across different fairness metrics, quantifying the trade-offs between accuracy and fairness through variation in algorithmic structure. Overall, our findings demonstrate the robustness of the original framework and its ability to transparently visualize algorithmic trade-offs structured under complex real-world conditions. These conditions are well illustrated through the highly regulated environment that the lending case study presents.
4.1. Fairness Incompatibility and Trade-off Analysis
Racial disparity measures such as those discussed in lee21, do not account for historical bias that is prevalent in mortgage data, although they are often used in evaluating lending decisions. For example, existing algorithmic fairness metrics either measure disparity in lending decisions at group or individual levels under absolute mathematical conditions as previously discussed in Section 2.2. Hence, we tested the compatibility of different mathematical notions of fairness (see Section 2 and Figure 2 in lee21) to decide if trade-offs such as those suggested in studies such as the aforementioned, Kleinberg2016, and hellman2020measuring are necessary for the models in this work.